# Non-perturbative approach for scalar particle production in Higgs-$R^{2}$ inflation

### v1

@@ Abstract
We investigate the non-perturbative production of scalar particles within the Higgs-$R^{2}$ inflation model, focusing on a massive spectator field that interacts gravitationally via a non-minimal coupling to the Ricci scalar $R$. By transforming the model to the Einstein frame, where direct couplings between inflationary dynamics and the spectator field emerge, we analyze both perturbative and non-perturbative particle production. We solve the spectator field equation numerically, using the Bogolyubov transformation, to calculate the comoving particle density. Our findings indicate that particle production occurs for light particles with conformal coupling and for masses $m_{\chi}>M/2$ in the minimally coupled scenario.
@@
Inflation, multi-field inflation, scalar fields, particle production
††preprint: APS/123-QED
I Introduction
@@
Cosmic inflation is the cornerstone of modern cosmology, providing a framework to understand the early universe and predict its fundamental characteristics. It describes a period of accelerated expansion driven by a scalar field $\phi$ slow-rolling towards the minimum of its potential $V(\phi)$. This mechanism addresses classical problems of the standard cosmological model and has become a central paradigm in cosmology [1, 2]. During the inflationary epoch, significant particle production occurs, especially as the kinetic energy of the inflaton field approaches its potential energy. This particle production can arise from gravitational effects due to the universe’s expansion and the dynamics of the inflaton field, serving as a classical background for propagating quantum fields [3, 4]. Known as gravitational particle production, this is a fundamental mechanism in quantum field theory in curved spacetime, offering insights into the origins of dark matter and the preheating and reheating epochs following inflation [5]. Several scenarios have been proposed for vacuum particle production during the early universe, particularly during inflation due to the evolution of the inflaton field and the universe’s expansion. Typically, the inflaton field is coupled to matter fields (bosonic, fermionic, or gauge fields) by a coupling constant $g$. At a critical value $\phi=\phi_{*}$ of the inflaton field, resonance begins and particle production is instantaneous [6, 7]. However, matter fields need not be directly coupled with the inflaton or standard model particles. Massive particles can be produced purely through gravitational effects, either via direct couplings with the gravitational field or induced couplings through a conformal transformation. If the matter field’s energy density contribution is negligible, the inflationary dynamics remain unaffected, leading to no back-reaction effects on spacetime geometry. Such fields are called spectator fields, and this mechanism provides a natural explanation for the origin of dark matter [8, 9, 10, 11, 12, 13]. Beyond dark matter origins, gravitational particle production is linked to significant early universe phenomena such as preheating after inflation [14, 15, 16, 17], primordial gravitational waves [18, 19, 20] and primordial magnetogenesis [21, 22, 23, 24]. Additionally, it imprints significant observable effects on the Cosmic Microwave Background (CMB), particularly in the generation of primordial non-Gaussianities [25, 26, 27].
In this work, we study gravitational particle production during and immediately after inflation driven by the SM Higgs boson and the scalaron field in the Higgs-$R^{2}$ inflation model. This model of inflation extends the SM Higgs inflationary model [28] to address the problem of large coupling [29, 30]. Here, cosmic inflation is driven by both the Higgs field and the extra scalar degree of freedom from the quadratic $R$ term, known as the scalaron. Unlike other multi-field inflation models, the Higgs-$R^{2}$ model effectively reduces to a single-field model during inflation, with the Higgs field acting as an auxiliary field while the scalaron drives inflation [31]. At the end of inflation during the reheating phase, both fields and the Ricci scalar oscillate around the potential’s minimum. This oscillatory behaviour enhance the scalar particle production through the non-perturbative mechanisms as tachyonic inestabilities [32, 33].
@@
We propose an extension to the Higgs-$R^{2}$ model by introducing a massive scalar field $\chi$ as a spectator field that it not contributing to the energy density during the inflationary epoch with negligible interactions with the SM but considering a direct coupling to Ricci scalar via a non-minimal coupling term $\xi_{\chi}R\,\chi^{2}$. Initially, there is no explicit coupling between the inflationary sector and the spectator field in the Jordan frame; however, gravitational interactions between the scalaron, the Higgs field and free scalars induce couplings when we change the conformal frame. This process results in gravitational particle production with significant implications for preheating after inflation and background dynamics. Induced coupling in the Einstein frame between free scalars and scalar sector of the action provides an excellent channel of scalaron decay into light scalars. However, in the non-perturbative approach the field mass $m_{\chi}$ does not necessarily have to be light and we can explore the possibility of scalar particle production for spectator field masses $m_{\chi}>M/2$ where $M\simeq 1.3\times 10^{-5}\,M_{p}$ is the scalaron mass. We take into account the oscillatory behavior of background dynamics after inflation as a classical background, which results in particle production due to the time-dependent background inducing a time-dependent frequency in the mode equations for the spectator field $\chi$. To achieve this, we use the Bogoliubov approach [3] to compute the comoving number density $n_{k}(t)$ of particles created by the background dynamics in Higgs-$R^{2}$ inflation by solving the mode equations for the quantized modes $\chi$.
@@
The article is organized as follows: section 2 provides a brief introduction to the Higgs-$R^{2}$ model and the background dynamics during and after inflation for a set of suitable initial conditions. Section 3 analyzes the couplings induced by gravitational effects in the Einstein frame, showing the decay ratio of scalaron into light free scalars. This result indicates that the perturbative approach is kinematically valid only for masses $m_{\chi}<M/2$ during the reheating epoch after inflation when $|R|<M$. In section 4 we derive the mode equation for the spectator field, which is necessary to investigate the behaviour of the time-dependent frequency during and after inflation. We adapt a numerical approach to solve the mode equation for modes that leave $5$ e-folds before the horizon crossing during inflation, which ensures that each mode is within the horizon at the start of the numerical simulation.
@@
Section 4 focuses on massive particle production during preheating. Finally, section 5 discusses the results and conclusions. In the following, we will use natural units in wihch ℏ$\hbar=1$, $c=1$, $\kappa_{\mathrm{B}}=1$ and we use the mass Planck as $8\pi G=M_{p}^{-2}$. We also work in the context of a spatially flat, homogeneous and isotropic universe described by the FLRW metric
$\mathrm{d}s^{2}$ $=$ −$-\mathrm{d}t^{2}+a^{2}(t)(\mathrm{d}x^{2}+\mathrm{d}y^{2}+\mathrm{d}z^{2})$
@@
II Brief review of Higgs-$R^{2}$ inflationary model
In this section, we will give a brief review of the Higgs-$R^{2}$ model in both Einstein and Jordan frames. For further details, we recommend the main references [31, 34, 30, 29] and the references therein. Let us get started the discussion in the Jordan frame; the model is giving by the action
$S_{\mathrm{J}}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}h^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}\right.\\[5.69046pt] \left.-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]\,,$ (2)
where $h$ is the Higgs field in the unitary gauge, the subindex $J$ denotes variables defined in the Jordan frame, $\xi_{h}$ is the constant coupling between Higgs field and the gravitational sector, $\xi_{s}$ is a mass parameter and $\lambda\sim 10^{-3}$ is the quartic self constant coupling of the Higgs field. This model contains a quadratic term in $R_{J}$ described by the constant $\xi_{s}$, wihch comes from the $R^{2}$ model [35] . Such constant is related to the scalaron mass as
$\xi_{s}=\dfrac{M_{p}^{2}}{3M^{2}}\,,$ (3)
where $M\simeq 1.3\times 10^{-5}M_{p}$ in order to be consistent with the value of the amplitude of primordial perturbations $A_{\zeta}$ [36]. The model (2) can be regarded as a UV extension of Higgs inflation, so it is interesting to make a double interpretation of the gravitational coupling in this conformal frame. In Higgs inflation, the constant coupling $\xi_{h}$ must fulfill the condition $\xi_{h}^{2}/\lambda\sim 10^{10}$ in order to be consistent with CMB observations [28, 37]. This causes that the cutoff scale of the model to be below the natural energy scale (Planck scale) during inflation, making the predictions of the model questionable; however, it has been shown that including the Ricci scalar cuadratic term modifies the value of the energy scale, rasinig it up to the Planck scale. To solve the strong copuling problem present in the scalar sector, the parameter $\xi_{s}$ has to obey the inequality [29]
@@ II Brief review of Higgs-$R^{2}$ inflationary model
This inequality allow us to make a perturbative treatment of the model. In order to describe the inflationary dynamics of the model, it is more convenient to bring the action (2) into the canonical Einstein-Hilbert form. So that we need to carry out a conformal transformation $\Omega^{2}(x)$ introducing an extra scalar degree of freedom $\phi$ called scalaron, which comes from the quadratic term in $R_{J}$. Thus, to go from Jordan frame to Eintein frame one has to perform a conformal transformation of the space-time metric $g_{\mu\nu}\to\Psi^{2}(x)g_{\mu\nu}$ with $\Psi(x)>0$, where the correct form of it is
@@
where $\alpha=\sqrt{2/3}M_{p}^{-1}$. The conformal transformation (5) removes the quadratic term of the scalar curvature $R_{J}$ introducing the scalaron $\phi$. Then in the Einstein frame the Higgs-$R^{2}$ model of inflation takes the form of a two-sacalar fields model with non-canonical kinetic term
$S_{E}=\int\mathrm{d}^{4}x\,\sqrt{-g}\,\left\{\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi\right.\\[5.69046pt] \left.-\dfrac{1}{2}\,e^{-\alpha\phi}g^{\mu\nu}\partial_{h}\,\partial_{\nu}h-V(\phi\,,h)\right\}\,,$ (6)
which can be wirtten in the general action for multi-field inflationary models
$S=\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\dfrac{M_{p}^{2}}{2}R\right.\\[5.69046pt] \left.-\dfrac{1}{2}G_{IJ}\,g^{\mu\nu}\partial_{\mu}\phi^{I}\,\partial_{\nu}\phi^{J}-V(\phi^{I})\right]\,,$ (7)
@@
We can see that the non-canonical kinetic term is associated with the geometry of the field space via the metric $G_{IJ}$ [38]. On the other hand, the potential in the Einstein frame has interaction terms between the Higgs field $h$ and the scalaron $\phi$, and takes the form
@@
Figure 1: Higgs-Starobinsky potential $V(\phi\,,h)/M_{p}^{4}$ in the Einstein frame (9). It is clear the position of the hill at $h=0$ and the position of the two valleys.
The potential $V(\phi^{I})$ depends on both the $\phi$ and $h$, and mixes terms which make more difficult analyze analytically the background trayectories. Nevertheless, during inflation we can make an effective approximation of the model such that inflation is driven by the single scalar field $\phi$ [31], and with the following interesting features
@@
    $V(\phi^{I})$ shows two valleys and one ridge
@@
    For $\xi_{h}$, $\xi_{s}$ possitive and enough large, the potential presents an attractor behavior
@@
    If $\phi\gg M_{p}$ with arbitrary initial conditions, the fields $\phi$ and $h$ will always evolve towards one of the valleys and thus beginning the slow-roll evolution
@@
    If $\phi<<M_{p}$, the potential is reduced to Higgs potential
@@
    The condition $\lambda_{\mathrm{eff}}\lesssim 4\pi$ is neccessary for perturbative preheating
@@
We can see the positions of the valleys and the hill at $h=0$ in the figure 1. The inflationary dynamics ocurr along one of these valleys, showing a universal attractor behavior, so that inevitably the fields $\phi$ and $h$ will always fall into one of the valleys and then drive slow-roll inflation. To find the valleys, we can perform the analysis in both Einstein [31] and Jordan [39] frames. In the Einstein frame in the limit $\phi\gg M_{p}$ (it is required for enough long inflation) and $h\gg v$, where $v=246\,\mathrm{GeV}$, the valleys are found out by the condition
∂$\dfrac{\partial V}{\partial h}=0\,,$ (13)
@@
Because of the potential (9) is symmetric respect to transformation $h\to-h$, the two valleys are localizated by the condition (14). Thus, this shows that the Higgs field is an auxiliary field during inflation, so if we put the result (14) back into the Einstein potential (9), we give the effective potential along the valleys
@@
Notice that the effective potential reduces to Higgs inflation if $\lambda\xi_{s}\ll\xi_{h}^{2}$ while one get $R^{2}$ inflation for $\lambda\xi_{s}\gg\xi_{h}^{2}$. In order to provide the correct amplitude of primordial scalar perturbations, the amplitude of (15) has to fulfill [40]
@@
where $N_{*}\sim 50-60$ is the number of e-folds before the end of inflation, and $A_{\zeta}\sim 2\times 10^{-9}$ is the amplitude of the primordial power spectrum. For $\lambda=0.013$ and $\xi_{h}\,,\xi_{s}>0$ which satisfy (16), the inflationary dynamics is effectively the same as single-field slow-roll inflation. In this paper we will only analyze the standard parameter region $\xi_{h}\,,\xi_{s}\,,\lambda>0$. Specifically, if we have $\xi_{h}\gg 0$ the isocurvature effects of the model are absent and the bending is whithin the weak mixing case $\eta_{\perp}\to 0$ [31], and the isocurvature and adiabatic modes are not coupled so that they evolve independetly in the standard parameter region of Higgs-$R^{2}$ model. Therefore, the description of the dynamics inside the valleys is an effective single-field inflation.
II.1 Background dynamics
What follows is to analyze the dynamics of the model. We assume that during inflation both scalaron $\phi$ and higgs field $h$ are homogeneous and only depend on cosmic time $t$. In this case the equations of motion can be written as
$\ddot{\phi}+3H\dot{\phi}+\dfrac{\partial V}{\partial\phi}$ $=$ −$-\dfrac{\alpha}{2}\,e^{-\alpha\phi}\dot{h}^{2}\,,$ (17)

### v2

@@ Abstract
Gravitational particle production in cosmology is a mechanism through which particles of different natures are produced during the very early universe. It is a general mechanism that explains how the universe became populated with the particles of the Standard Model after cosmic inflation and may also account for the origin of dark matter. In this work, we study the non-perturbative production of massive scalar particles in the Higgs-$R^{2}$ inflation model, a two-field scalar inflation model within the Einstein frame. We consider spectator scalar fields that are conformally or minimally coupled to the gravitational field through the curvature scalar $R$ which in turn is a time-dependent function determined by the fields driving inflationary dynamics. We numerically compute the production of these particles using the Bogolyubov transformation method for each scenario, aiming to assess the spectrum of the produced particles. For both scenarios, we consider light particles with masses $m_{\chi}\ll H_{\mathrm{end}}$ and large masses that exceed the Hubble scale at the end of inflation $m_{\chi}\gtrsim H_{\mathrm{end}}$.
keywords:
@@
Particle production, Multi-field inflation, Higgs inflation, Early universe
1 Introduction
Cosmic inflation is the cornerstone of modern cosmology, providing a framework to understand the early universe and predict its fundamental characteristics.
@@ 1 Introduction
It describes a period of accelerated expansion driven by a scalar field $\phi$ that slowly rolls towards the minimum of its potential $V(\phi)$. This mechanism addresses classical problems of the standard cosmological model and has become a central paradigm in cosmology [1, 2]. During the inflationary epoch, significant particle production occurs, particularly when the kinetic energy of the inflaton field becomes comparable to its potential energy. This particle production can arise from gravitational effects due to the universe’s expansion and the dynamics of the inflaton field, serving as a classical background for propagating quantum fields [3, 4]. This phenomenon, known as gravitational particle production, is a fundamental mechanism in quantum field theory in curved spacetime, offering insights into the origins of dark matter and the preheating and reheating epochs following inflation [5]. Several scenarios have been proposed for vacuum particle production during the early universe, particularly during inflation due to the evolution of the inflaton field and the universe’s expansion. Typically, the inflaton field is coupled to matter fields (bosonic, fermionic, or gauge fields) by a coupling constant $g$. At a critical value $\phi=\phi_{*}$ of the inflaton field, resonance is triggered, leading to instantaneous particle production [6, 7]. However, matter fields need not be directly coupled with the inflaton or standard model particles. Massive particles can be produced purely through gravitational effects, either via direct couplings with the gravitational field or induced couplings through a conformal transformation. If the contribution of the matter field to the energy density is negligible, the inflationary dynamics remain unaffected, leading to no back-reaction effects on spacetime geometry. Such fields are called spectator fields, and this mechanism provides a natural explanation for the origin of dark matter [8, 9, 10, 11, 12, 13]. Beyond its role in the origin of dark matter, gravitational particle production is also linked to key early-universe phenomena such as preheating after inflation [14, 15, 16, 17], primordial gravitational waves [18, 19, 20], and primordial magnetogenesis [21, 22, 23, 24]. Furthermore, it leaves significant observable imprints on the Cosmic Microwave Background (CMB), particularly in the generation of primordial non-Gaussianities [25, 26, 27].
@@
In this work, we study gravitational particle production during and immediately after inflation driven by the SM Higgs boson and the scalaron field in the Higgs-$R^{2}$ inflation model. This inflation model extends the SM Higgs inflationary model [28] to address the problem of large coupling [29, 30]. Here, cosmic inflation is driven by the Higgs field and the extra scalar degree of freedom from the quadratic $R$ term, known as the scalaron. Unlike other multi-field inflation models, the Higgs-$R^{2}$ model effectively reduces to a single-field model during inflation, with the Higgs field acting as an auxiliary field while the scalaron drives inflation [31]. At the end of inflation during the reheating phase, both fields and the Ricci scalar oscillate around the potential minimum. This oscillatory behavior enhances scalar particle production through non-perturbative mechanisms such as tachyonic instabilities [32, 33]. We propose an extension to the Higgs-$R^{2}$ model by introducing a massive scalar field $\chi$ as a spectator field that does not contribute to the energy density during the inflationary epoch, with negligible interactions with the SM but considering a direct coupling to the Ricci scalar via a non-minimal coupling term $\xi_{\chi}R\,\chi^{2}$. Initially, the inflationary sector and the spectator field are not explicitly coupled in the Jordan frame; however, gravitational interactions between the scalaron, the Higgs field, and free scalars induce couplings upon transforming to the Einstein frame. This process results in gravitational particle production with significant implications for preheating after inflation and background dynamics. Induced coupling in the Einstein frame between free scalars and the scalar sector of the action provides an excellent channel for scalaron decay into light scalars. However, in the non-perturbative approach, the field mass $m_{\chi}$ does not necessarily have to be light, and we can explore the possibility of scalar particle production for spectator field masses $m_{\chi}\gtrsim H_{\mathrm{end}}$, where $H_{\mathrm{end}}$ is the Hubble ratio at the end of inflation. We take into account the oscillatory behavior of background dynamics after inflation as a classical background, which results in particle production due to the time-dependent background inducing a time-dependent frequency in the mode equations for the spectator field $\chi$. To achieve this, we employ the Bogoliubov approach [3] to compute the comoving number density $n_{k}(t)$ of particles generated by the background dynamics in Higgs-$R^{2}$ inflation, solving the mode equations for the quantized modes $\chi$.
@@
The article is organized as follows: Section 2 provides a brief introduction to the Higgs-$R^{2}$ model and the background dynamics during and after inflation for a set of suitable initial conditions. Section 3 analyzes the couplings induced by gravitational effects in the Einstein frame, in Section 4, we derive the mode equation for the spectator field, which is necessary to investigate the behavior of the time-dependent frequency during and after inflation. We adopt a numerical approach to solve the mode equation for modes that leave 5 e-folds before the horizon crossing during inflation, ensuring that each mode is within the horizon at the start of the numerical simulation. Section 4 examines the production of massive particles during preheating. Finally, Section 5 discusses the results and conclusions.
@@
In the following, we will use natural units in which $\hbar=1$, $c=1$, $\kappa_{\mathrm{B}}=1$ and we use the mass Planck as $8\pi G=M_{p}^{-2}$. We also work in the context of a spatially flat, homogeneous, and isotropic universe described by the FLRW metric
$\mathrm{d}s^{2}$ $=$ $-\mathrm{d}t^{2}+a^{2}(t)(\mathrm{d}x^{2}+\mathrm{d}y^{2}+\mathrm{d}z^{2})$
@@
2 Brief review of Higgs-$R^{2}$ inflationary model
In this section, we will give a brief review of the Higgs-$R^{2}$ model in both Einstein and Jordan frames. For further details, we recommend the main references [31, 34, 30, 29] and the references therein. Let us get started with the discussion in the Jordan frame; the model is given by the action
$S_{\mathrm{J}}=\int\mathrm{d}x^{4}\sqrt{-g_{J}}\left[\dfrac{1}{2}(M_{p}^{2}+\xi_{h}h^{2})R_{J}+\dfrac{\xi_{s}}{4}\,R_{J}^{2}-\dfrac{1}{2}g^{\mu\nu}_{J}\,\partial_{\mu}h\,\partial_{\nu}h-\dfrac{\lambda}{4}\,h^{4}\right]\,,$ (2)
where $h$ is the Higgs field in the unitary gauge, the subindex $J$ denotes variables defined in the Jordan frame, $\xi_{h}$ is the constant coupling between the Higgs field and the gravitational sector, $\xi_{s}$ is a mass parameter and $\lambda$ is the quartic self constant coupling of the Higgs field. This model contains a quadratic term in $R_{J}$ described by the constant $\xi_{s}$, which comes from the $R^{2}$ model [35]. Such a constant is related to the scalaron mass as $\xi_{s}=M_{p}^{2}/3M^{2}$. The model (2) can be regarded as a UV extension of Higgs inflation, so it is interesting to make a double interpretation of the gravitational coupling in this conformal frame. In Higgs inflation, the constant coupling $\xi_{h}$ must fulfill the condition $\xi_{h}^{2}/\lambda\sim 10^{10}$ to be consistent with CMB observations [28, 36]. This causes the cutoff scale of the model to be below the natural energy scale (Planck scale) during inflation, making the predictions of the model questionable; however, it has been shown that including the Ricci scalar quadratic term modifies the value of the energy scale, raising it to the Planck scale. To solve the strong coupling problem present in the scalar sector, the parameter $\xi_{s}$ has to obey the relation [29]
@@ 2 Brief review of Higgs-$R^{2}$ inflationary model
This expression allows us to make a perturbative treatment of the model. To describe the inflationary dynamics of the model, it is more convenient to bring the action (2) into the canonical Einstein-Hilbert form. So we need to carry out a conformal transformation $\Omega^{2}(x)$ introducing an extra scalar degree of freedom $\phi$ called scalaron, which comes from the quadratic term in $R_{J}$. Thus, to go from the Jordan frame to Einstein frame one has to perform a conformal transformation of the space-time metric $g_{\mu\nu}\to\Psi^{2}(x)g_{\mu\nu}$ with $\Psi(x)>0$, where the correct form of it is
@@
where $\alpha=\sqrt{2/3}M_{p}^{-1}$. The conformal transformation (4) removes the quadratic term of the scalar curvature $R_{J}$ introducing the scalaron $\phi$. Then in the Einstein frame, the Higgs-$R^{2}$ model of inflation takes the form of a two-scalar fields model with a non-canonical kinetic term
$S_{E}=\int\mathrm{d}^{4}x\,\sqrt{-g}\,\left\{\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}g^{\mu\nu}\partial_{\mu}\phi\,\partial_{\nu}\phi-\dfrac{1}{2}\,e^{-\alpha\phi}g^{\mu\nu}\partial_{h}\,\partial_{\nu}h-V(\phi\,,h)\right\}\,,$ (5)
which can be written in the general action for multi-field inflationary models
Figure 1: Scalar potential in the Einstein frame of the Higgs-$R^{2}$ model seen from different angles. The curve $\phi=\mathbf{cte}$ has the shape of the potential of the $R^{2}$ inflation model, while $h=\mathrm{cte}$ has the shape of the quartic Higgs potential of the Standard Model. We have restricted ourselves to the standard parameter space $\xi_{h}=4000$, $\xi_{s}=2\times 10^{9}$ and $\lambda=0.013$.
$S=\int\mathrm{d}^{4}x\,\sqrt{-g}\left[\dfrac{M_{p}^{2}}{2}R-\dfrac{1}{2}G_{IJ}\,g^{\mu\nu}\partial_{\mu}\phi^{I}\,\partial_{\nu}\phi^{J}-V(\phi^{I})\right]\,,$ (6)
@@
We can see that the non-canonical kinetic term is associated with the geometry of the field space via the metric $G_{IJ}$ [37]. On the other hand, the potential in the Einstein frame has interaction terms between the Higgs field $h$ and the scalaron $\phi$ and takes the form
@@
The potential $V(\phi^{I})$ depends on both the $\phi$ and $h$ and mixes terms which make it more difficult to analyze analytically the background trajectory. Nevertheless, during inflation, we can make an effective approximation of the model such that inflation is driven by the single scalar field $\phi$ [31] and with the following interesting features
@@
    $V(\phi^{I})$ shows two valleys and one ridge.
@@
    For $\xi_{h}$, $\xi_{s}$ positive and enough large, the potential presents an attractor behavior.
@@
    If $\phi\gg M_{p}$ with arbitrary initial conditions, the fields $\phi$ and $h$ will always evolve towards one of the valleys thus beginning the slow-roll evolution.
@@
    If $\phi\ll M_{p}$, the potential is reduced to Higgs potential
@@
    The condition $\lambda_{\mathrm{eff}}\lesssim 4\pi$ is necessary for perturbative preheating
@@
We can see the positions of the valleys and the hill at $h=0$ in figure 1. The inflationary dynamics occur along one of these valleys, showing a universal attractor behavior, so that inevitably the fields $\phi$ and $h$ will always fall into one of the valleys and then drive slow-roll inflation. To find the valleys, we can perform the analysis in both Einstein [31] and Jordan [38] frames. In the Einstein frame in the limit $\phi\gg M_{p}$ (it is required for enough long inflation) and $h\gg v$, where $v=246\,\mathrm{GeV}$, the valleys are found out by the condition
$\dfrac{\partial V}{\partial h}=0\,,$ (12)
@@
Because the potential (8) is symmetric with respect to transformation $h\to-h$, the two valleys are localized by the condition (13). Thus, this shows that the Higgs field is an auxiliary field during inflation, so if we put the result (13) back into the Einstein potential (8), we give the effective potential along the valleys
@@
Notice that the effective potential reduces to Higgs inflation if $\lambda\xi_{s}\ll\xi_{h}^{2}$ while one gets $R^{2}$ inflation for $\lambda\xi_{s}\gg\xi_{h}^{2}$. To provide the correct amplitude of primordial scalar perturbations, the amplitude of (14) has to fulfill [39]
@@
where $N_{*}\sim 50-60$ is the number of e-folds before the end of inflation, and $A_{\zeta}\sim 2\times 10^{-9}$ is the amplitude of the primordial power spectrum. For $\lambda=0.013$ and $\xi_{h}\,,\xi_{s}>0$ which satisfy (15), the inflationary dynamics is effectively the same as single-field slow-roll inflation. In this paper, we will only analyze the standard parameter region $\xi_{h}\,,\xi_{s}\,,\lambda>0$. Specifically, if we have $\xi_{h}\gg 0$ the isocurvature effects of the model are absent the bending is within the weak mixing case $\eta_{\perp}\to 0$ [31], and the isocurvature and adiabatic modes are not coupled so that they evolve independently in the standard parameter region of Higgs-$R^{2}$ model. Therefore, the description of the dynamics inside the valleys is an effective single-field inflation.
2.1 Background dynamics
What follows is to analyze the dynamics of the model. We assume that during inflation both scalaron $\phi$ and Higgs field $h$ are homogeneous and only depend on cosmic time $t$. In this case, the equations of motion can be written as
$\ddot{\phi}+3H\dot{\phi}+\dfrac{\partial V}{\partial\phi}$ $=$ $-\dfrac{\alpha}{2}\,e^{-\alpha\phi}\dot{h}^{2}\,,$ (16)